### v2

# DIVERSE: A Dataset of YouTube Video Comment Stances with a Data Programming Model

Abstract
Stance detection of social media text is a key component of many real-world applications like evaluating marketing campaigns, evaluating political policies or candidates, or evaluating information environments. However, creating automatic stance labeling systems requires the manual annotation of stances, which is both tedious and resource-intensive. This paper introduces an stance labeling method that makes use of weak signals of sentence tone, then consolidating these signals with a Data Programmingmodel for the final stance label. In a time of international conflict, understanding the public opinion towards the country’s military is crucial for recruitment. We present DIVERSE, a dataset involve stances towards YouTube videos of the US military [^1] Datasetavailableathttps://doi.org/10.5281/zenodo.10493803. On average, the videos have 200 comments each, and the stances skew slightly towards the “against” characterization for both the US army and the video.
Index Terms:
stance detection, social media, large language models, ensemble learning
I Introduction
Today’s world is rife with armed conflict: the Russia-Ukraine war, the Israel-Hamas conflict, the Yemeni Civil War, Somali Civil War… and so forth. The US is involved in a number of these conflicts, and also has the mammoth task of defending her own. During conflict, it is crucial to know the stance of the people towards its own defending forces, as a gauge of the support the forces have from the public. However, US military is facing a recruitment crisis. For the last few years, the US Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023). Social media plays an important role in discerning stances (AlDayel and Magdy, 2021), and in recruitment (Ng and Cruickshank, 2023). Thus, it is critical to understand how messages from the military is received on social media, for it may be used to shape beliefs of users. More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, be it a war recruiting effort or a organizational recruitment effort or for influencing consumer behavior.
A particular indicator of the reception of social media content is stance. Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social studies. Several articles have proposed methods and datasets for the task of stance detection (Allaway and McKeown, 2023; Du Bois, 2007). However, existing datasets explore stance as expressed on a single social media platform, mainly X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020). Since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), along with varying definitions of stance (Ng and Carley, 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels. This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
To address these shortcomings in prior stance dataset, we propose a new stance dataset based on YouTube comments and labeled with machine labeling through the data programming paradigm of weak supervision. This dataset is a collection of YouTube comments posted on videos uploaded by the US Army [^2] Offical US Army YouTube Channel: https://www.youtube.com/@usarmy. The comments are annotated for weak labels representing tone (hate speech, sarcasm, sentiment), inference of two LLMs (stance towards video, stance towards army), then combining these labels to produce the final stance labels of the comment towards the US Army. With this dataset, we develop a automated and systematic approach to labeling stances, aiding in understanding critical issues with the public opinion towards the army and its recruitment efforts.
Our contributions of this paper are as follows:
  • The use of an data-programming based stance-labeling methodology. We combine recent advances for creating large, high-quality labeled datasets Ratner et al. (2016, 2019) with contextual knowledge to generate quality stance labels. This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
  • A novel benchmark stance dataset of YouTube comments. We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets. While most stance datasets use posts from Twitter, ours provides posts from the YouTube platform instead, contributing to the breadth of datasets. Another difference is that our dataset labels stances towards more than one target in a comment (the only other being Sobhani et al. (2017)), and provides labels to weak signals including: the presence of contentious issues, conspiracy theories, and other emotional content. Finally, this dataset has more than three times the number of labeled comments over the current largest benchmark dataset (Conforti et al. (2020) at ${}32,000$ data points vs 173,013 data points).
II Related Work
Stance Datasets
The task of stance detection has led to the development of several datasets. Ng and Carley (2022) consolidated several stance detection datasets with social media posts across multiple topics. These datasets use three basic labels – favor, against, neither – to express how the sentence views a target (Mohammad et al., 2016). There are datasets that use labels of implicit and explicit support and denial to indicate the overtness of stance expression (Qazvinian et al., 2011). Automated stance detection algorithms have been constructed with these datasets using support vector machines (Elfardy and Diab, 2016), logistic regressions (Augenstein et al., 2016), neural network models (Fang et al., 2019; Siddiqua et al., 2019) and deep learning models (Küçük and Can, 2020). Recently, Large-Language Models have also been used for identifying stance in sentences (Zhang et al., 2023a; Gatto et al., 2023).
YouTube Datasets
To date, there have been several research investigations involving YouTube, but few datasets. For example, there are a number of papers investigating influence campaigns (Marcoux et al., 2021; Hussain et al., 2018), conspiracy theories (Liaw et al., 2023), comment toxicity (Obadimu et al., 2019), and even cross-platform use of YouTube videos (Ginossar et al., 2022). Despite the focus on analyzing YouTube data, including comment data, most research projects do not release their datasets (Liaw et al., 2023). Furthermore, to the best of the authors’ knowledge, there are no datasets or investigations of YouTube data for stance detection, particularly concerning the comments section on YouTube videos. It is important to note, however, that there are a number of commercially available social media suites to analyze comments to posts on platforms for indicators like sentiment and influencer engagement, but none of these tools provide stance detection (Ng and Carley, 2022; Rogan et al., 2022).
Large Language Models for Stance Detection
Recently, researchers have explored the application of Large Language Models (LLMs) in stance detection. Existing studies primarily focused on ChatGPT, yielding mixed outcomes. Zhang et al. (2022) demonstrated improved results on the SemEval2016 benchmark dataset using ChatGPT with instruction-based prompts. Other works have also found good, if not necessarily state-of-the-art performance, on the task of stance detection for LLMs (Ziems et al., 2023; Liyanage et al., 2023; Zhang et al., 2023b). Conversely, Aiyappa et al. (2023) raised concerns about potential data contamination from ChatGPT’s extensive training dataset, casting doubt on the reliability of evaluations. Mets et al. (2023) examined ChatGPT’s usability as a zero-shot classifier for stance detection, reporting competitive but ultimately inferior performance compared to supervised models. Despite these findings, the effectiveness of LLMs in stance classification remains promising, if uncertain, especially without fine-tuning on labeled data or prompt engineering.
Weak Supervision for Dataset creation
Weak Supervision has emerged as a promising approach for expeditiously generating large training datasets. This methodology leverages heuristic rules and weak supervision to programmatically label data at scale (Ratner et al., 2016, 2019). Rather than relying solely on manually annotated datasets, weak supervision enables the automatic creation of training sets by combining various weak sources, derived from heuristics, distant supervision, and domain expertise. These weaker, noisy labels can then be combined to produce high-quality labels. This paradigm shift offers an efficient means of tackling data scarcity challenges, especially in scenarios where obtaining labeled data is time-consuming or expensive. The incorporation of weak supervision into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
III Dataset Creation
This dataset contains comments of YouTube comments, each labeled for a set of two final stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army. There are three possible options for stances: “supports”, “against”, and “neutral”. This dataset also contains the weak labels used to fold into the final stance through weak supervision.
III-A Data Collection
This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military. We collected comments on the videos of the US Army’s official YouTube channel (username: @usarmy [^3] https://www.youtube.com/usarmy) using the YouTube Data API [^4] https://developers.google.com/youtube/v3/docs/comments/list. This returned both comments as well as replies to comments. Collection took place from 2 to 5 October 2023, where we retrieved a total of 173,013 comments across 858 videos. For each comment, we collected the following information:
  • id: the unique id of a comment
  • comment: the text of the comment
  • author: a unique id for the author of the comment
  • like_count: the likes the comment received
  • published_at: the time and date of the comment’s publication
  • conversation_id: a unique id of the conversation that a comment belongs to. This is also the id of the root comment
  • video_id: the unique id of the video the comment is commenting on
  • name: the title of the video the comment is commenting on
  • description: a text description of the video the comment is commenting on
  • timestamp: the time and date of the publication of the video the comment is commenting on
III-B Data Annotation Methodology
Instead of using human annotators for stance labeling, we used weak supervision Ratner et al. (2016, 2019). This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled. Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension Liu and Jaidka (2023) and negotiation strategies using the presence of reasoning and friendliness Jaidka et al. (2023). We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection. These labels were then combined through a weak supervision model to produce the final labels. Figure 1, depicts the entire data annotation methodology workflow. We detail key aspects of the workflow in the following sections.
Figure 1: Pipeline of stance labeling methodology note: change LLM box to 2 boxes not 3
III-C Annotating Weak Labels
To label the comments for stance, we first produced weak labels. To do this we leveraged dataset-specific traits, domain expertise and LLMs. For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army. Therefore, the presence of these weak signals can be combined to indicate the final label. It is important to note for any given weak label, it need not produce a label for every comment in the dataset. Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
III-C1 Hate Speech
Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion Fortuna and Nunes (2018). In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the US Army more generally featured aspects of hate speech, like antisemitic and misogynistic content. Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection Fortuna and Nunes (2018). For this weak label, we used the dehatebert model from HuggingFace [^5] Hate-speech-CNERG/dehatebert-mono-english Aluru et al. (2020), a multi-lingual fine-tuned BERT model. This model was used as is without any fine-tuning. For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the “hate” response as an “against” stance label.
III-C2 Sarcasm
Sarcasm is a literary device to convey one’s intention in an implicit manner Chaudhari and Chandankhede (2017). As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video. We used a multilingual sarcasm detector666helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian. For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
III-C3 Sentiment
Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic Ibrohim et al. (2023). In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment. We used a multilingual-cased sentiment analysis model777lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model. For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”. We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
III-C4 Use of Keywords
We manually determined a set of keywords and phrases that upon present, would represent either one of the stance labels. For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label. For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance.
III-C5 Stance from Large Language Models
Based on the recent work demonstrating effectiveness of LLMs in the stance classification task (Ng et al., 2024), we also directly labeled stances using LLMs.
Pre-tuning of LLMs
Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware. Specifically, we used two different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture and (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture. We used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023). For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020). Based on the research in (Ng and Carley, 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets. For the training of the Mistral LLM, we used a LoRA adapter (Hu et al., 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term. The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter, 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps. To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works. The following was the training prompt:
“‘Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following social media statement and determine its stance towards the provided [dataset entity].
Respond with a single word: “for”, “against”,“neutral”.
[dataset entity]: {entity}
statement: {statement to classify}
stance: ”’
Where the [dataset entity] was varied depending on the benchmark dataset (e.g., for Election2016, it was ‘politician’, while for phemerumors it was ‘rumor’). For the other variation of the prompt, the “Analyze…” sentence was replaced with a question-based tasking in the form of: “What is the stance of the following social media statement towards the provided [dataset entity]”. For the srq dataset, we changed the prompt slightly to incorporate the post that the reply is replying to as the task in that dataset is to label the stance of the reply to the post, versus the stance of the post about a topic or person. The loss was determined by whether the model could output the tokens for the correct stance, and only those tokens. The training of the model was done on a local computer running Ubuntu Linux with a 96-core processor and three A6000 GPUs. Overall, this resulted in three, different LLMs for labeling the comments: one off-the-shelf UL2 model and two, fine-tuned Mistral models.
Analyzing stances with LLMs
In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt. The labeling prompt without incorporating the reply structure is:
“‘Analyze the following YouTube comment to a video posted by the U.S. Army named “{title}” and determine its stance towards the provided entity. Respond with a single word: “for”, “against”, “neutral”, or “unrelated”. Only return the stance as a single word, and no other text.
entity: {entity}
comment: {comment to classify}
stance:
while the labeling prompt that did incorporate the reply structure is:
“‘Analyze the following YouTube comment to a video posted by the U.S. Army named “{title}” and determine its stance towards the provided entity. If the statement is a reply to another YouTube comment, that YouTube comment is listed below. Respond with a single word: “for”, “against”, “neutral”, or “unrelated”. Only return the stance as a single word, and no other text.
entity: {entity}
previous comment: {previous_comment}
comment: {comment to classify}
stance:
As with the training prompt, the variations for each prompt came from changing the task part of the prompt (i.e., the “Analyze…”) sentence to a question-based task prompt. These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt. Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
III-C6 Use of Analogous Stances
Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label. We used the stance of the comment toward the US Army. Generally speaking, if a comment was against the US Army, it would also be against a video posted by the US Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al. (2017); Zhang et al. (2023b). Thus, we used the final stance toward the US Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video. It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the US Army but would be very against a particular video.
III-D Final Stance Determination
Having created the weak labels for the comments, we then use the weak supervision methodology in the form of the snorkel Python package [^8] https://snorkel.readthedocs.io/en/master/index.html, which uses a unified generative model that takes into account partial coverage of the weak labels as well as correlations between the weak labels to produce the final labels. We refer the reader to Ratner et al. (2016) and Ratner et al. (2019) for the specifics on the function, how it is trained, and how it produces the final labels. For the final labels, we allow the stance to take the form of “supports”, “against”, and “neutral” for both the stance toward the video and the stance toward the US Army.
III-E Human evaluation
After machine annotation, we had four individuals annotate how many? what percentage of total? comments. For each comment, the authors independently annotated the stance towards the US Army and the stance towards the video. The manual annotations were then analyzed for inter-annotator agreement, and the machine annotations were compared against the manual ones.
IV Data Properties and Analysis
This section describes some exploratory data analysis performed on this dataset. The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection) on the US Army’s official YouTube Channel.
IV-A Dataset Properties
Distribution of videos over time
Generally speaking, the channel has published an increasing number of videos over time, with some occasional periods of fewer videos. Figure 3 depicts the distribution of all the videos on the channel over time. Table II summarizes the statistics for the comments to each video. Most videos within the data set only have a single comment, with the number of comments per video being highly skewed and some videos having a large number of comments (the most commented-on video, ‘Soldier wakes up and chooses “HOOAH!”’, has 9.4% of all of the comments by itself). When it comes to the timing of the comments themselves, we find the overwhelming have occurred recently in the channel’s timeline (since 2022).
Figure 2: Distribution of comments over time
From Figure 2, the plot of comments over time, there are significant spikes in comments in August, September, and October of 2022 and again in May and June of 2023. In the case of the first spike, the most commented-on video was ‘Soldier wakes up and chooses “HOOAH!”’, while the second spike in the Spring of 2023 has ‘Words of wisdom from a #WWII veteran’ the most commented-on video. For the first video, it seems the commentary is divisively split about what the Army does or should do in the world, while in the second case the comments seem to both positively reminisce about the Army in World War II and lament about its current state. Turning to the comments themselves, we find that most comments are relatively short, like many social media posts, but a few comments are much longer at over 1,000 words for the comment. Table I has the word count statistics for the comments.
Word-level Statistics of Comments Value
Mean 16.77
Standard Deviation 32.67
Minimum 1.00
25th Percentile 4.00
50th Percentile (Median) 9.00
75th Percentile 18.00
Maximum 1851.00
TABLE I: Word-level Statistics of the Comments
Statistics of comments per video Value
Mean 201.65
Standard Deviation 873.19
Minimum 1
25th Percentile 5
50th Percentile (Median) 27
75th Percentile 77
Maximum 16289
Mode 1
TABLE II: Per Video Comment Statistics
Figure 3: Distribution of videos published over time
IV-B Stance Label Properties
Turning now to stance labels produced through our methodology for the dataset, we find that the labels skew slightly toward the “against” characterization for both the US Army and the videos being posted on the channel. Table III summarizes the label counts for each of the stances.
Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label. In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like Hossain et al. (2020) and Kochkina et al. (2018), the “neutral” stance tends to be the majority stance label. Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior. Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
Stance U.S. Army Video
against 63676 (34.87%) 67405 (36.92%)
neutral 49153 (26.91%) 49310 (27.01%)
supports 60183 (32.99%) 56297 (30.88%)
TABLE III: Labels Counts (and percentages of the total number) for Each of the Stance Targets
Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time. Figure 4 displays the distribution of the different classes of stance toward the US Army for all of the comments. The different stances toward the videos follow a very similar trend. We further observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels. In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”. Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against” or “neutral”. Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
Figure 4: Distribution of videos published on the U.S. Army’s Official Channel over time
To better understand these differences in stances, we looked at the videos with the most positive, neutral and negative receptions. Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video. On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the US Army. The video had 66.9% of comments that were against it. And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
Finally, since this dataset was labeled with more than one stance, we can also look at the relationships between the two stances. Figure 5, displays a confusion matrix between the stance toward the U.S. Army and the stance toward the videos. Not unsurprisingly, we find these stances closely resemble each other. It is, however, worth pointing out that we do see a significant number of comments that are supportive of the video, but neutral toward the Army as well as against or neutral to the video, but supportive of the Army. In the first case, from some manual inspection, these are often comments that give a positive word about the video and nothing else. For example, a comment like “that video was awesome!” fits into this category. For, the second category these comments are often comments that express some frustration or anger at a particular video, but are otherwise supportive of the Army. For example, a comment like “What is this bs? The Army is better than this…” Overall, the two stance labels closely resemble each other in this dataset, but there are some notable exceptions where they differ, which highlights some of the complexity of stance within this domain and more generally.
Figure 5: Confusion Matrix of Stance Toward the U.S. Army and Stance Toward the Video
IV-C Comparison with human evaluation
The inter-annotator agreement was Iain, results here. Comparing the machine-based annotations with the human-based annotations, we get an accuracy of how many. elaborate on accuracy
IV-D Leave-One-Out Stance Detection
To further validate the stance labels produced by the proposed weak supervision method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley, 2022). We trained a model on the stance labels and then evaluated that model on the other benchmark datasets. In our test, we trained a cased BERT model [^9] https://huggingface.co/bert-base-cased (Devlin et al., 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets. The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels. The text that was input to the model was given as:
“target: {target of stance} [SEP] comment: {comment}”
In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in Ng and Carley (2022). Table IV has the out-of-domain test results. The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043. This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers Ng and Carley (2022). This therefore suggests that the stance labels proposed in this paper are valid stance labels.
Dataset Macro F1 Score
semeval2016 0.3517
election2016 0.3439
covid-lies 0.3020
phemerumors 0.2295
wtwt 0.3018
srq 0.2971
Average 0.3043
TABLE IV: Macro F1 Scores for Model Applied to Different Datasets
V Discussion
Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join the forces. A positive opinion facilitates army recruitment and support, especially in times of crisis. Social media engagement plays a pivotal role in shaping this opinion. This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
To construct this dataset, we employed weak supervision, which leverages weak labels for automated labeling. This paradigm provides a transition from direct human annotation to machine-assisted annotation, allowing scalability to larger datasets. For our specific dataset, we utilized subject matter expertise, domain knowledge, and recent advancements in using LLMs for stance determination to create weak labels. Specifically, we used weak labels such as the presence of hate speech, sarcasm, and sentiment as indicators for certain stances. This transformation converted the challenging, ambiguously defined problem of determining and labeling stance into a more well-defined problem for specific aspects or stances. These could then be aggregated through weak supervision to produce comprehensive stance labeling. For instance, the presence of hate speech indicates a comment against the video and, by extension, against the user who uploaded it (e.g., the Army).
We further enhanced these weak labels with contributions from LLMs. Currently, the effectiveness of LLMs in tasks like stance labeling remains uncertain. While recent research shows promise for using LLMs, especially when combined with prompt engineering, for stance detection and labeling, the generalizability of these methods and their compatibility with all LLMs is unclear. Through weak supervision, we treat these LLM-derived labels as noisy yet valuable, enhancing the creation of a stance-labeled dataset. The combination of dataset-specific weak signals and the linguistic understanding power of LLMs produces high-quality labels for stance in a more scalable manner compared to purely manual or machine-generated labels.
Finally, our Leave-One-Out experiment produces a standard macro F1 score of around 0.34, a score that is similar to past work. This suggests that the stance labels produced by the proposed methodology are indeed valid. This result also highlights findings from past work that trained stance detection models do not necessarily generalize to other data domains (e.g., YouTube versus Twitter, or targets of videos versus political figures, etc.). Thus, we suggest training a stance classification model using multiple benchmark datasets so the model can better recognize the nuances of the stances and targets involved. The low F1 score also indicates the difficulty of the stance classification task, whereby language models do struggle to differentiate the nuances in opinions towards a target entity, despite having helper annotations like hate and sentiment to provide indications of stances. This, therefore, opens up avenues for exploration of improving the stance classification methods, especially on out-of-domain tasks, to provide a more accurate label towards an entity and therefore more accurate downstream insights.
Potential Applications
The dataset proposed in this work has several potential uses. First, this dataset presents a versatile resource for research on the analysis of stance in YouTube video comments. Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks. Notably, it originates from YouTube, offering a departure from the prevalent X-based datasets. Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances). The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.
Given the dataset’s proximity to a politically and culturally divisive topic — the military — it encompasses a spectrum of online behaviors, including the propagation of conspiracy theories, trolling, hate speech, and misinformation. As such, it serves as a valuable resource for studies on misinformation and disinformation.
Finally, the dataset’s temporal collection allows researchers to explore dynamic aspects of stance-taking behavior over time. This temporal dimension provides insights into the evolution of discussions and sentiments within the context of the US Army’s YouTube channel. In summary, the dataset offers a rich foundation for various studies, ranging from advancing stance detection methodologies to delving into the nuanced dynamics of online interactions related to a prominent and controversial subject.
Dataset Availability
We collected data provided by the YouTube API and did not perform additional scraping. We only collected public videos and comments, and made no attempt to collect private information. We anonymized the author IDs and did not perform any author-level analysis. The code used for construction of the dataset and the dataset itself is provided at https://doi.org/10.5281/zenodo.10493803.
Limitations and Future Work
Our dataset is limited to the comments from one channel, and the comments we could collect (not all videos posted to the channel allowed comments). We hope to expand our dataset to include the comments towards videos posted by armies around the world, to evaluate public opinion towards the army, across the world. In addition, our dataset is limited to relatively small LLMs due to computational constraints. In the future, we hope to expand the dataset annotation to different types of LLMs with varying prompting schemes to better understand the stance annotation capabilities of LLMs.
VI Conclusion
Understanding the stance expressed in YouTube comments towards an entity is essential for gaining insights into public opinion regarding the authors of the videos. Here, we introduce the DIVERSE dataset, a dataset that captures the stances towards the US Army. This dataset is formulated from comments gathered from the US Army’s YouTube videos. We constructed the final stance variable utilizing weak labels, including the identification of hate speech and sarcasm, and incorporated stance inferences from Large Language Models. By employing weak supervision, we amalgamated these diverse labels to generate the final stance annotations. We anticipate that this dataset will prove invaluable to researchers investigating public sentiments and stances toward military entities.
Acknowledgements
The research for this paper was supported in part by the Center for Informed Democracy and Social-cybersecurity (IDeaS) at Carnegie Mellon University. This work was also conducted within the Cognitive Security Research Lab at the Army Cyber Institute. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Defense, the U.S. Army, or the U.S. Government.